# Lagrangian chaos and unique ergodicity for stochastic primitive equations

### v1

@@ 7. Approximate controllability and weak irreducibility
subject to the initial conditions $\mathbf{x}_{0}=\mathbf{x}\in\mathbb{T}^{3}$, $\xi_{0}=\xi\in\mathbb{S}^{2}$, $\widecheck{\xi}_{0}=\widecheck{\xi}\in\mathbb{S}^{2}$ and $\mathcal{A}_{0}=\mathrm{Id}_{{\mathbb{R}}^{3\times 3}}$. In these equations, $\Pi_{\xi}=\mathrm{Id}-(\xi\otimes\xi)$ denotes the orthogonal projection from ${\mathbb{R}}^{3}$ onto the tangent space of $\mathbb{S}^{2}$ at the point $\xi$, where $\xi$ is viewed as an element of ${\mathbb{R}}^{3}$ with unit length. Additionally, $\top$ represents the matrix transpose.
@@
The proof extends the arguments in [16, Lemma 7.1] to the current context. The main idea is roughly as follows. First we prove that during the time interval $[0,1/2]$ we can move the point $\mathbf{x}_{t}$ from $\mathbf{a}$ to $\mathbf{a}^{\prime}$ while $\xi_{t}$ is moved to some $\xi^{\prime}\in\mathbb{S}^{2}$, and in a second step we prove that during the time interval $[1/2,1]$ we can move $\xi_{t}$ from $\xi^{\prime}$ to $\eta^{\prime}$ without changing the position of $\mathbf{x}_{t}$. Moreover, the construction shows also that $v_{0}=v_{1}=0$. The proof naturally splits into two steps.
Step 1: There exists a smooth control $Qg$ such that support in time of $Qg$ is contained in $(0,1/2)$ and a unit vector $\xi^{\prime}\in\mathbb{S}^{2}$ such that
@@
Note that solutions to the 2D Navier-Stokes equations also solve the primitive equations (by considering $z$ as a dummy variable). Therefore, in light of the proof of [16, Lemma 7.1], we can assume that during the interval $(0,1/4)$, we have moved the position of the horizontal variables. In particular, it is enough to show the claim of Step 1 whenever $\Pi_{x,y}\mathbf{a}=\Pi_{x,y}\mathbf{a}^{\prime},$ where $\Pi_{x,y}$ is the projection onto the first two components and $(0,1/2)$ is replaced by $(1/4,1/2)$. Let $\mathbf{a}=(a,b,c)$ and $\mathbf{a}^{\prime}=(a,b,c^{\prime})$ for some $a,b\in\mathbb{T}$ and $c,c^{\prime}\in(0,1)$ (here we used that $\mathbf{a},\mathbf{a}^{\prime}\in\mathcal{O}$). We can assume that either $c<c^{\prime}$ or $c^{\prime}<c$; otherwise, there is nothing to prove. Without loss of generality, we may assume $c<c^{\prime}$; the other case is analogous. Let $f:\mathbb{T}\to{\mathbb{R}}$ be a mean free smooth function, which will be fixed later. Furthermore let us fix $\varphi\in C_{{\rm c}}((1/4,1/2))$ such that
@@
Since $f$ is mean-free, we have $\int_{0}^{1}\nabla_{x,y}\cdot v_{1,t}(\cdot,\zeta)\,\mathrm{d}\zeta=0$.
Set $F(z)\stackrel{{\scriptstyle{\rm def}}}{{=}}\int_{0}^{z}f(\zeta)\,\mathrm{d}\zeta$ and note that
$[w(v_{1,t})](\mathbf{x})=-\int_{0}^{z}(\nabla_{x,y}\cdot v_{1,t})(x,y,\zeta)\,\mathrm{d}\zeta=\varphi_{t}\cos(y-b)F(z).$
Hence, the induced vector field $u$ is given by
@@
Hence $y_{t}\equiv b$ and the last equation simplifies to
$\dot{z}_{t}=\varphi_{t}F(z_{t}),\qquad z_{1/4}=c.$
Next, we choose a smooth mean-free map $f$ for which the unique solution to the former ODE satisfies $z_{1/2}=c^{\prime}$. Recall that we are assuming $10\leq c<c^{\prime}<$ 1. Construct $f$ such that $f|_{[0,c^{\prime}]}=1$ and $f$ is a mean-free smooth periodic map (this can be achieved as $10\leq c<c^{\prime}<$ 1). Note that $F|_{[0,c^{\prime}]}=z$ and therefore the unique solution of the above ODE is given by $z_{t}=c\exp(\int_{1/4}^{t}\varphi_{s}\,\mathrm{d}s)$ for all $t$ such that $z_{t}<1$. Since $\varphi_{t}$ satisfies (7.5), we have $z_{1/2}=c^{\prime}$. Step 1 now follows by taking:
@@
Recall that $Q^{-1}:\mathbb{H}^{r+\alpha}_{0}\to\mathbb{H}^{r}_{0}$ for all $r>0$ by Assumption 2.2. Therefore the above formula is well-defined due to the regularity of $(v_{1,t})_{t}$ and $\int_{\mathbb{T}^{3}}v_{1,t}\,\mathrm{d}x=0$ and $\int_{\mathbb{T}^{3}}\mathbb{P}[(v_{1,t}\cdot\nabla_{x,y})v_{1,t}+w(v_{1,t})\partial_{z}v_{1,t}]\,\mathrm{d}x=0$ as $\nabla\cdot u_{1,t}=0$ by construction.
Step 2: Let $\xi^{\prime}$ be as in Step 1. There exists a smooth control $Qg_{2}$ such that support in time of $Qg_{2}$ is contained in $[1/2,3/4]$ and
@@
i.e. the control changes the direction $\xi_{t}$ from $\xi$ to $\xi^{\prime}$ while keeping fixed the position $\mathbf{a}$. As in Step 1, due to the proof of [16, Lemma 7.1], it is enough to show that we can produce flows which allow us to rotate the direction $\xi_{t}$ around the $z$-axis during the time interval $(1/2,3/4)$. Hence, let us fix $\theta\in(0,2\pi)$ and pick $\psi\in C_{{\rm c}}((1/2,3/4))$ such that
@@
The proof of the existence of such an $f$ is elementary. Let $F(\zeta)\stackrel{{\scriptstyle{\rm def}}}{{=}}\int_{0}^{z}f(z)\,\mathrm{d}z$. Arguing as in Step 1, for $\mathbf{x}=(x,y,z)\in\mathbb{T}^{3}$,
@@
Finally, let us note that for all $n\in\mathbb{S}^{2}$ we have $An\perp n$ and therefore
@@
The proof of (7.13) follows from [16, Proposition 7.4] as solutions to the 2D Navier-Stokes equations are $z$-independent solutions to the 3D primitive equations. In other words, we consider solution $v_{t}$ to (7.9) that are independent of $z\in(0,1)$. For simplicity, we take $t=1$. Fix $\psi\in C^{\infty}_{{\rm c}}(0,t)$ such that $\int_{0}^{1}\psi_{r}\,\mathrm{d}r=\log(1+M)$ and let $\mathbf{a}=(a,b,c)$. To prove (7.13) one takes for all $t\in[0,1]$ and $\mathbf{x}=(x,y,z)\in\mathbb{T}^{3}$
@@
(7.10) $\mathbf{P}\big{(}(v_{T},\mathbf{x}_{T})\in B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathbb{T}^{3}}(\mathbf{a}^{\prime},\varepsilon)\,|\,(v_{0},x_{0})=(v,\mathbf{a})\big{)}>0,$
@@
where $K_{\sigma,\varepsilon}(\mathbf{a}^{\prime},\eta^{\prime})=B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathbb{T}^{3}}(\mathbf{a}^{\prime},\varepsilon)\times B_{\mathbb{S}^{2}}(\eta^{\prime},\varepsilon)$. Finally, for all $M>0$, $\mathbf{a}\in\mathbb{T}^{3}$ and $(t,\varepsilon)$ as above,
(7.13) $\mathbf{P}\big{(}(v_{T},\mathbf{x}_{T},\mathcal{A}_{T})\in B_{\varepsilon}(0)\times B_{\varepsilon}(\mathbf{a})\times B_{M}^{\complement}(0)\,|\,(0,\mathbf{x},\mathcal{A}_{0})=(0,\mathbf{a},\mathrm{Id})\big{)}>0.$
@@
Here, we partially follow the proof in [16, Proposition 2.15, pp. 1976]. We only prove the assertion for the Lagrangian process; the other claim follows similarly. Let $\widetilde{\mu}$ be an arbitrary invariant measure for $(u_{t},\mathbf{x}_{t})$ on $\mathbb{H}^{\sigma}_{0}\times\mathbb{T}^{3}$. By Proposition 3.2, one can check that $\widetilde{\mu}$ is concentrated on $\mathbb{H}^{\sigma^{\prime}}_{0}\times\mathbb{T}^{3}$ for all $\sigma^{\prime}<\sigma+1$, where $\sigma$ as in (2.14). Fix $\sigma^{\prime}\in(\sigma,\sigma+1)$ and let $N\gg 1$ be such that
$\widetilde{\mu}(\mathcal{S})>1/2\qquad\text{ and }\qquad\mathcal{S}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{\|v\|_{\mathbb{H}^{\sigma^{\prime}}}\leq N\}\times\mathbb{T}^{3}$
Next, let us collect some facts about the (deterministic) PEs without forcing, i.e.,
@@
(7.15) $\mathbf{P}((v_{T_{\delta}},\mathbf{x}_{T_{\delta}})\in B_{\mathbb{H}^{\sigma}}(0,\delta)\times B_{\mathbb{T}^{3}}(\mathbf{x},\delta)\,|\,(v_{0},\mathbf{x}_{0})=(v,\mathbf{x}))>0.$
In the following, if no confusion seems likely, we write $B(0,\delta)$ and $B(\mathbf{x},\delta)$ instead of $B_{\mathbb{H}^{\sigma}}(0,\delta)$ and $B_{\mathbb{T}^{3}}(\mathbf{x},\delta)$, respectively.
Now, fix $\varepsilon>0$, $(v,\mathbf{x})\in\mathcal{S}$ and $\mathbf{x}^{\prime}\in\mathcal{O}$. Let $\varepsilon^{\prime}>0$ be as in Lemma 7.3 for $T=1$. Recall that $L_{t}$ denotes the Markov semigroup associated with the Lagrangian process $(v_{t},\mathbf{x}_{t})$. By the semigroup property, it follows that
$\mathbf{P}((v_{T_{\varepsilon^{\prime}}+1},\mathbf{x}_{T_{\varepsilon^{\prime}}+1})\in B(0,\delta)\times B(\mathbf{x}^{\prime},\delta)\,|\,(v_{0},x_{0})=(v,\mathbf{x}))$
$=[L_{T_{\delta}}(L_{1}(\mathbf{1}_{B(0,\delta)\times B(\mathbf{x}^{\prime},\delta)}))](v,\mathbf{x})$
$=\int_{\Omega}\mathbf{P}\big{(}(v_{1},\mathbf{x}_{1})\in B(0,\delta)\times B(\mathbf{x}^{\prime},\delta)\,|\,v_{0}=v_{T_{\varepsilon^{\prime}}}(\widetilde{\omega}),\mathbf{x}_{0}=\mathbf{x}_{T_{\varepsilon^{\prime}}}(\widetilde{\omega})\big{)}\,\mathrm{d}\mathbf{P}(\widetilde{\omega})\stackrel{{\scriptstyle(i)}}{{>}}0.$
where $v_{T_{\delta}}(\widetilde{\omega})$ and $\mathbf{x}_{T_{\delta}}(\widetilde{\omega})$ are the solution to the stochastic PEs (1.1) with initial data $(v,\mathbf{x})$ for the noise realization $\widetilde{\omega}\in\Omega$ and the associated solution to the Lagrangian flow (7.1). Moreover, in $(i)$ we used (7.15) as well as (7.10).
@@
$\widetilde{\mu}(B(0,\delta)\times B(\mathbf{x}^{\prime},\delta))\geq\int_{\mathcal{S}}L_{T_{\delta^{\prime}+1}}((v,\mathbf{x}),\,B(0,\delta)\times B(\mathbf{x}^{\prime},\delta))\,\mathrm{d}\widetilde{\mu}(v,\mathbf{x})>0.$
@@
Hence, the arbitrariness of $\delta$ and $\mathbf{x}^{\prime}\in\mathcal{O}$ implies $\mathrm{supp}\,\,\widetilde{\mu}\supseteq\overline{\{0\}\times\mathcal{O}}=\{0\}\times\mathbb{T}^{3}$. ∎
Acknowledgments

### v2

@@ 7. Approximate controllability and weak irreducibility
subject to the initial conditions $\mathbf{x}_{0}=\mathbf{x}\in\mathcal{O}$, $\xi_{0}=\xi\in\mathbb{S}^{2}$, $\widecheck{\xi}_{0}=\widecheck{\xi}\in\mathbb{S}^{2}$ and $\mathcal{A}_{0}=\mathrm{Id}_{{\mathbb{R}}^{3\times 3}}$. In these equations, $\Pi_{\xi}=\mathrm{Id}-(\xi\otimes\xi)$ denotes the orthogonal projection from ${\mathbb{R}}^{3}$ onto the tangent space of $\mathbb{S}^{2}$ at the point $\xi$, where $\xi$ is viewed as an element of ${\mathbb{R}}^{3}$ with unit length. Additionally, $\top$ represents the matrix transpose.
@@
The proof extends the “shear flows” method in [16, Lemma 7.1] to the context of PEs. The main idea is roughly as follows. First we prove that during the time interval $[0,1/2]$ we can move the point $\mathbf{x}_{t}$ from $\mathbf{a}$ to $\mathbf{a}^{\prime}$ while $\xi_{t}$ is moved to some $\xi^{\prime}\in\mathbb{S}^{2}$, and in a second step we prove that during the time interval $[1/2,1]$ we can move $\xi_{t}$ from $\xi^{\prime}$ to $\eta^{\prime}$ without changing the position of $\mathbf{x}_{t}$. Moreover, the construction also shows that $v_{0}=v_{1}=0$. The proof naturally splits into two steps.
Step 1: There exists a smooth control $g$ such that support in time of $g$ is contained in $(0,1/2)$ and a unit vector $\xi^{\prime}\in\mathbb{S}^{2}$ such that
@@
Note that solutions to the 2D Navier-Stokes equations also solve the PEs (by considering $z$ as a dummy variable). Therefore, in light of the proof of [16, Lemma 7.1], we can assume that during the interval $(0,1/4)$, we have moved the position of the horizontal variables. In particular, it is enough to show the claim of Step 1 whenever $\Pi_{x,y}\mathbf{a}=\Pi_{x,y}\mathbf{a}^{\prime},$ where $\Pi_{x,y}$ is the projection onto the first two components and $(0,1/2)$ is replaced by $(1/4,1/2)$. Let $\mathbf{a}=(a,b,c)$ and $\mathbf{a}^{\prime}=(a,b,c^{\prime})$ for some $a,b\in\mathbb{T}$ and $c,c^{\prime}\in(0,1)$. We can assume that either $c<c^{\prime}$ or $c^{\prime}<c$; otherwise, there is nothing to prove. Without loss of generality, we may assume $c<c^{\prime}$; the other case is analogous. Let $f:\mathbb{T}\to{\mathbb{R}}$ be a mean-zero smooth function, which will be fixed later. Furthermore let us fix $\varphi\in C_{{\rm c}}((1/4,1/2))$ such that
@@
Since $f$ is mean zero, we have $\int_{0}^{1}\nabla_{x,y}\cdot v_{1,t}(\cdot,z)\,\mathrm{d}z=0$ on $\mathbb{T}^{2}_{x,y}$.
Set $F(z)\stackrel{{\scriptstyle{\rm def}}}{{=}}\int_{0}^{z}f(z^{\prime})\,\mathrm{d}z^{\prime}$ and note that
$[w(v_{1,t})](\mathbf{x})=-\int_{0}^{z}(\nabla_{x,y}\cdot v_{1,t})(x,y,z^{\prime})\,\mathrm{d}z^{\prime}=\varphi_{t}\cos(y-b)F(z).$
Hence, the induced vector field $u_{1,t}$ is given by
@@
From the above, it also follows that $y_{t}\equiv b$, and the last equation simplifies to
(7.6) $\dot{z}_{t}=\varphi_{t}F(z_{t}),\qquad z_{1/4}=c.$
Next, we choose a smooth mean-zero map $f$ for which the unique solution to the former ODE satisfies $z_{1/2}=c^{\prime}$. To this end, recall that we are assuming $10<c<c^{\prime}<$ 1. Let $f$ be such that $f|_{[0,c^{\prime}]}=1$ and $f$ is a mean zero smooth periodic map. The existence of such $f$ is evident because $10\leq c<c^{\prime}<$ 1. In particular, $F|_{[0,c^{\prime}]}=z$ and the unique solution of the ODE (7.6) is given by
$z_{t}=c\exp(\int_{1/4}^{t}\varphi_{s}\,\mathrm{d}s)$
for all $t$ such that $z_{t}<1$. Since $\varphi_{t}$ satisfies (7.5), we have $z_{1/2}=c^{\prime}$. The claim of Step 1 now follows by taking:
@@
Indeed, since $Q^{-1}:\mathbb{H}^{r+\alpha}_{0}\to\mathbb{H}^{r}_{0}$ for all $r>0$ by Assumption 2.2, the above formula is well-defined due to the regularity of $(v_{1,t})_{t}$, the mean zero of $v_{1,t}$ and $\int_{\mathbb{T}^{3}}\mathbb{P}[(v_{1,t}\cdot\nabla_{x,y})v_{1,t}+w(v_{1,t})\partial_{z}v_{1,t}]\,\mathrm{d}x=0$ as $\nabla\cdot u_{1,t}=0$.
Step 2: Let $\xi^{\prime}$ be as in Step 1. There exists a smooth control $g_{2}$ such that support in time of $g_{2}$ is contained in $[1/2,3/4]$ and
@@
i.e. the control changes the direction $\xi_{t}$ from $\xi$ to $\xi^{\prime}$ while keeping fixed the position $\mathbf{a}$. As in Step 1, following the proof of [16, Lemma 7.1], it suffices to demonstrate that we can construct flows that enable us to rotate the direction $\xi_{t}$ around the $x$- and $y$-axis during the time interval $(1/2,3/4)$. We provide the details only for the $x$-axis, as the case for the $y$-axis is analogous.
Fix $\theta\in(0,2\pi)$ and pick $\psi\in C_{{\rm c}}((1/2,3/4))$ such that
@@
The proof of the existence of such an $f$ is elementary (indeed, it is enough to take a smooth function $f$ satisfying the last two pointwise conditions in (7.9), and then modify appropriately the mapping far from the point $c^{\prime}\in(0,1)$). Let $F(\zeta)\stackrel{{\scriptstyle{\rm def}}}{{=}}\int_{0}^{z}f(z)\,\mathrm{d}z$. Arguing as in Step 1, for $\mathbf{x}=(x,y,z)\in\mathbb{T}^{3}$,
@@
Finally, let us note that for all $n\in\mathbb{S}^{2}$ we have $\Gamma n=(0,-n_{z},n_{y})\perp n$ and therefore
@@
The proof of (7.14) follows from [16, Proposition 7.4] as solutions to the 2D Navier-Stokes equations are $z$-independent solutions to the 3D primitive equations. In other words, we consider solution $v_{t}$ to (7.10) that are independent of $z\in(0,1)$. For simplicity, we take $t=1$. Fix $\psi\in C^{\infty}_{{\rm c}}(0,t)$ such that $\int_{0}^{1}\psi_{r}\,\mathrm{d}r=\log(1+M)$ and let $\mathbf{a}=(a,b,c)$. To prove (7.14) one takes for all $t\in[0,1]$ and $\mathbf{x}=(x,y,z)\in\mathbb{T}^{3}$,
@@
(7.11) $\mathbf{P}\big{(}(v_{T},\mathbf{x}_{T})\in B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathcal{O}}(\mathbf{a}^{\prime},\varepsilon)\,|\,(v_{0},x_{0})=(v,\mathbf{a})\big{)}>0,$
@@
where $K_{\sigma,\varepsilon}(\mathbf{a}^{\prime},\eta^{\prime})=B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathcal{O}}(\mathbf{a}^{\prime},\varepsilon)\times B_{\mathbb{S}^{2}}(\eta^{\prime},\varepsilon)$. Finally, for all $M>0$, $\mathbf{a}\in\mathbb{T}^{3}$ and $(t,\varepsilon)$ as above and $J_{\sigma,\varepsilon,M}=B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathcal{O}}(\mathbf{a},\varepsilon)\times(B_{{\mathbb{R}}^{3\times 3}}(0,M))^{\complement}$,
(7.14) $\mathbf{P}\big{(}(v_{T},\mathbf{x}_{T},\mathcal{A}_{T})\in J_{\sigma,\varepsilon,M}\,|\,(0,\mathbf{x},\mathcal{A}_{0})=(0,\mathbf{a},\mathrm{Id})\big{)}>0.$
@@
Here, we partially follow the proof in [16, Proposition 2.15, pp. 1976]. We only prove the assertion for the Lagrangian process; the other claim follows similarly. Let $\widetilde{\mu}$ be an arbitrary invariant measure for $(u_{t},\mathbf{x}_{t})$ on $\mathbb{H}^{\sigma}_{0}\times\mathcal{O}$. By Proposition 3.2, one can check that $\widetilde{\mu}$ is concentrated on $\mathbb{H}^{\sigma^{\prime}}_{0}\times\mathcal{O}$ for all $\sigma^{\prime}<\sigma+1$, where $\sigma$ as in (2.14). Fix $\sigma^{\prime}\in(\sigma,\sigma+1)$ and let $N\gg 1$ be such that
$\widetilde{\mu}(\mathcal{S})>1/2\qquad\text{ and }\qquad\mathcal{S}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{\|v\|_{\mathbb{H}^{\sigma^{\prime}}}\leq N\}\times\mathcal{O}$
Next, we collect some facts about the (deterministic) PEs without forcing, i.e.,
@@
(7.16) $\mathbf{P}((v_{T_{\delta}},\mathbf{x}_{T_{\delta}})\in B_{\mathbb{H}^{\sigma}}(0,\delta)\times B_{\mathcal{O}}(\mathbf{x},\delta)\,|\,(v_{0},\mathbf{x}_{0})=(v,\mathbf{x}))>0.$
In the following, if no confusion seems likely, we write $B(0,\delta)$ and $B(\mathbf{x},\delta)$ instead of $B_{\mathbb{H}^{\sigma}}(0,\delta)$ and $B_{\mathcal{O}}(\mathbf{x},\delta)$, respectively.
Now, fix $\varepsilon>0$, $(v,\mathbf{x})\in\mathcal{S}$ and $\mathbf{x}^{\prime}\in\mathcal{O}$. Let $\varepsilon^{\prime}>0$ be as in Lemma 7.3 for $T=1$. Moreover, let $T_{\varepsilon^{\prime}}$ be such that (7.16) holds with $\delta=\varepsilon^{\prime}$. Finally, recall that $L_{t}$ denotes the Markov semigroup associated with the Lagrangian process $(v_{t},\mathbf{x}_{t})$. By the semigroup property, it follows that
$\mathbf{P}((v_{T_{\varepsilon^{\prime}}+1},\mathbf{x}_{T_{\varepsilon^{\prime}}+1})\in B(0,\varepsilon)\times B(\mathbf{x}^{\prime},\varepsilon)\,|\,(v_{0},x_{0})=(v,\mathbf{x}))$
$=[L_{T_{\delta}}(L_{1}(\mathbf{1}_{B(0,\varepsilon)\times B(\mathbf{x}^{\prime},\varepsilon)}))](v,\mathbf{x})$
$\textstyle=\int_{\Omega}\mathbf{P}\big{(}(v_{1},\mathbf{x}_{1})\in B(0,\varepsilon)\times B(\mathbf{x}^{\prime},\varepsilon)\,|\,v_{0}=v_{T_{\varepsilon^{\prime}}}(\widetilde{\omega}),\mathbf{x}_{0}=\mathbf{x}_{T_{\varepsilon^{\prime}}}(\widetilde{\omega})\big{)}\,\mathrm{d}\mathbf{P}(\widetilde{\omega})\stackrel{{\scriptstyle(i)}}{{>}}0.$
where $v_{T_{\varepsilon^{\prime}}}(\widetilde{\omega})$ and $\mathbf{x}_{T_{\varepsilon^{\prime}}}(\widetilde{\omega})$ are the solution to the stochastic PEs (1.1) with initial data $(v,\mathbf{x})$ for the noise realization $\widetilde{\omega}\in\Omega$ and the associated solution to the Lagrangian flow (7.1). Moreover, in $(i)$ we used (7.16) as well as (7.11).
@@
$\widetilde{\mu}(B(0,\varepsilon)\times B(\mathbf{x}^{\prime},\varepsilon))\geq\int_{\mathcal{S}}L_{T_{\varepsilon^{\prime}+1}}((v,\mathbf{x}),\,B(0,\varepsilon)\times B(\mathbf{x}^{\prime},\varepsilon))\,\mathrm{d}\widetilde{\mu}(v,\mathbf{x})>0.$
@@
Hence, the arbitrariness of $\varepsilon$ and $\mathbf{x}^{\prime}\in\mathcal{O}$ implies $\mathrm{supp}\,\,\widetilde{\mu}\supseteq\{0\}\times\mathcal{O}$. ∎
Appendix A Some useful results on Feller Markov semigroups
Here, we gathered some results on the general theory of Markov processes, which were used in the proof of our main results in Subsection 3.3. The reader is referred to the text before Proposition 3.11 for the definition of the strong Feller property.
Lemma A.1.
Let $Z$ and $W$ be Polish spaces. Let $(z_{t})_{t}$ and $(z_{t},w_{t})_{t}$ be Markov processes on $Z$ and $Z\times W$, respectively. Assume that $(z_{t},w_{t})_{t}$ is strongly Feller on $Z\times W$. Let $\nu$ be an ergodic measure for $(z_{t})_{t}$. Suppose there exists a measure $\mu$ on $W$ with connected support and such that $\nu\times\mu$ is invariant for $(z_{t},w_{t})_{t}$. Then $\nu\otimes\mu$ is ergodic.
As the case analysed in Subsection 3.3, the above can be useful in case the uniqueness of invariant measures (and therefore ergodicity) cannot be easily proven via the Krein-Milman and a variant of the Khasminskii’s theorems (see [25, Proposition 4.1.1] for the latter).
The above result immediately implies
Corollary A.2.
Let $(z_{t})_{t}$ be a strongly Feller Markov process on a Polish space $Z$. Let $\nu$ be an invariant measure for $(z_{t})_{t}$ on $Z$ such that $\mathrm{supp}\,\nu$ is connected. Then $\mu$ is ergodic.
The connectedness assumption cannot be omitted in the above results. Indeed, in the situation of Corollary A.2, if a Feller Markov process $(z_{t})_{t}$ has two dinstict ergodic measures, then $\mu=\frac{1}{2}(\mu_{1}+\mu_{2})$ is invariant but not ergodic. However, if the strong Feller property holds, it has disconnected support as $\mathrm{supp}\,\mu_{1}\cap\mathrm{supp}\,\mu_{2}=\varnothing$.
Proof of Lemma A.1.
By a well-known characterisation of ergodic measures (see e.g., [25, Theorem 3.2.4]), it is sufficient to show that if a set $E\subseteq Z$ satisfies
(A.1) $P_{t}\mathbf{1}_{E}(z,w)=\mathbf{1}_{E}(z,w)$
for all $t>0$ and $\nu\otimes\mu$-almost all $(z,w)$, then either $\nu\otimes\mu(E)=0$ or $\nu\otimes\mu(E)=1$. In the previous, $(P_{t})_{t}$ denotes the semigroup associated to the Markov semigroup $(z_{t},w_{t})_{t}$, i.e., $P_{t}\mathbf{1}_{E}(z,w)=\mathbf{E}[\mathbf{1}_{E}(z_{t},w_{t})]$.
Next, we assume that (A.1) holds. Let $F_{0}\subseteq Z\times W$ be a full measure set on which the equality (A.1) holds for $t=1$. Clearly, $F_{0}$ is dense in $\mathrm{supp}\,\nu\times\mathrm{supp}\,\mu$. By the strong Feller and (A.1) with $t=1$, the mapping $F_{0}\ni(z,w)\mapsto P_{1}\mathbf{1}_{E}(z,w)=\mathbf{1}_{E}(z,w)$ can be extended to a continuous mapping on $\mathrm{supp}\,\nu\times\mathrm{supp}\,\mu$; which we denoted by $\psi_{E}$. From the equality (A.1) with $t=1$, the mapping $\psi_{E}$ takes values in $\{0,1\}$. In particular, the sets $A_{j}=\psi_{E}^{-1}(\{j\})$ for $j\in\{0,1\}$ are open and closed in $\mathrm{supp}\,\nu\times\mathrm{supp}\,\mu$, $A_{0}\cup A_{1}=\mathrm{supp}\,\nu\times\mathrm{supp}\,\mu$ and
(A.2) $\nu\otimes\mu(A_{1}\Delta E)=0\ \ \text{ where }\ \ A_{1}\Delta E\stackrel{{\scriptstyle{\rm def}}}{{=}}A_{1}\setminus E\cup E\setminus A_{1}.$
By connectedness of $\mathrm{supp}\,\mu$, it follows that $A_{j}=\widetilde{A}_{j}\times\mathrm{supp}\,\mu$ where $\widetilde{A}_{j}\in\mathscr{B}(Z)$. Indeed, if for instance, $A_{0}$ has not such a product structure, then there exist $z\in Z$ and $w,w^{\prime}\in W$ such that $(z,w)\in A_{0}$ and $(z,w^{\prime})\in A_{1}$. Since $A_{0},A_{1}$ are open and closed in $\mathrm{supp}\,\nu\times\mathrm{supp}\,\mu$, it follows that the sections $A_{j}^{z}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{w\in\mathrm{supp}\,\mu\,:\,(z,w)\in A_{j}\}$ are also open and closed. By connectedness of $\mathrm{supp}\,\mu$, either $A_{0}^{z}=\mathrm{supp}\,\mu$ or $A_{0}^{z}=\varnothing$; this contradicts the fact that $A_{0}$ does not have a product structure.
Hence, we proved that $A_{j}=\widetilde{A}_{j}\times\mathrm{supp}\,\mu$ for some $\widetilde{A}_{j}\in\mathscr{B}(Z)$. Due to (A.2) and the invariance of $\nu\otimes\mu$, it follows that $P_{t}\mathbf{1}_{A_{1}\Delta E}=0$, $\nu\otimes\mu$-almost everywhere. Therefore, (A.1) implies $\widetilde{P}_{t}\mathbf{1}_{\widetilde{A}_{1}}=\mathbf{1}_{\widetilde{A}_{1}}$ for all $t>0$ and $\nu$-almost everywhere. Here, $(\widetilde{P}_{t})_{t}$ denotes the Markov semigroup induced by $(z_{t})_{t}$. The ergodicity of $\nu$ implies either $\nu(\widetilde{A}_{1})=0$ or $\nu(\widetilde{A}_{1})=1$; and therefore either $\nu\otimes\mu(E)=0$ or $\nu\otimes\mu(E)=1$ due to (A.2), as desired. ∎
Finally, we prove a uniqueness result of invariant measures with the same marginals. With a slight abuse of notation, for a measure $\mu$ on the product space $Z\times W$, we denote by $\mu|_{Z}$ the marginal of $\mu$ on $Z$, i.e., $\mu|_{Z}(A)=\mu(A\times W)$ for all $A\in\mathscr{B}(Z)$.
Lemma A.3.
Let $Z$ and $W$ be Polish spaces. Let $(z_{t})_{t}$ and $(z_{t},w_{t})_{t}$ be Markov processes on $Z$ and $Z\times W$, respectively. Let $\nu$ be an ergodic measure for the Markov process $(z_{t})_{t}$, and let $\mu$ be an extreme point of the convex set
$\mathscr{Q}=\{\mu\text{ is an invariant measure for $(z_{t},w_{t})_{t}$ such that $\mu|_{Z}=\nu$}\}.$
Then $\mu$ is ergodic.
By the Krein-Milman theorem, the extremality of $\mu$ in $\mathscr{Q}$ follows trivially in case $W$ is compact. This situation appears in our application of Lemma A.3 where $W=\mathrm{Pr}_{2}$ is the two-dimensional projective space, see Subsection 3.3.
Proof.
Here, we employ the characterisation of ergodic measures as the extreme point of invariant measures, see e.g., [25, Proposition 3.2.7]. By contradiction, assume that $\mu$ is not ergodic. Hence, there exist $t\in(0,1)$ and distinct invariant measures $\mu_{1}$ and $\mu_{2}$ of $(z_{t},w_{t})_{t}$ such that $\mu=(1-t)\mu_{1}+t\mu_{2}.$ It is clear that the marginals measures $\mu_{1}|_{Z}$ and $\mu_{2}|_{Z}$ are invariant for $(z_{t})_{t}$. Moreover, since $\mu|_{Z}=\nu$, the marginals measures $\mu_{1}|_{Z}$ and $\mu_{2}|_{Z}$ satisfy $\nu=(1-t)\mu_{1}|_{Z}+t\mu_{2}|_{Z}$. The ergodicity of $\nu$ implies $\nu=\mu_{1}|_{Z}=\mu_{2}|_{Z}$ ([25, Proposition 3.2.7]). Hence, $\mu_{1},\mu_{2}\in\mathscr{Q}$ and the assumption $\mu=t\mu_{1}+(1-t)\mu_{2}$ with $\mu_{1}\neq\mu_{2}$ and $t\in(0,1)$ contradicts the extremality of $\mu$ in $\mathscr{Q}$. The proof is therefore complete. ∎
Acknowledgments.